# Redirection for Erasing Memory (REM): Towards a universal unlearning method for corrupted data

### v1

@@ 7 Discussion and conclusion
We presented a taxonomy for corrupted data unlearning tasks along two dimensions: discovery rate and regularity, and showed that no prior algorithm succeeds across the entire space. We then proposed REM, the first unlearning method that performs strongly across this 2D space of unlearning tasks. REM redirects corrupted data to a dedicated model part that is dropped or deactivated after unlearning to remove the influence of corrupted data from the model. A limitation of REM is that, as in ETD, its masks are binary. Future work could consider softer masking strategies that may allow for better self-organization within $\theta_{o_{2}}$, leading to masks of (even undiscovered) corrupted data having greater overlap with one another. Such better masking strategies may lead to bridging the performance gap with REM (IDEAL) in Fig 3. However, in its current version, REM already makes important strides forward as the first universal method that is strong across our task space. We believe that our 2D framework is an important step forward in better understanding when different families of unlearning algorithms fail, offering useful tools and vocabulary for explaining their behaviours. We hope that future work expands our framework through identifying other key dimensions, builds upon it, for instance by translating its tasks to other modalities, and leverages it for investigating the behaviours of algorithms for other types of unlearning applications, such as unlearning for privacy.
@@ Appendix A Appendix
A.1 Healing Metric
@@
A.2 Model and UL Method Parameters
We perform unlearning with a limited hyperparameter sweep as the best values are already known for high and lower regularity tasks and varying discovery rates from [9] (poison trigger and interclass confusion tasks). We report the selected hyperparameters in Tab. 3. The ViT architecture and setup for SVHN is adapted from [28], only changing the optimizer to Adam and adapting the learning rate and batch size accordingly. The ResNet-9 setup as well as the poison trigger and interclass confusion tasks are used from [9]. The random label unlearning problem is adapted from [18]. The REM parameter $\beta$ is kept at 1 to avoid adding extra complexity to the method by introducing weighting.. Training data splits, choice of which sample IDs to manipulate etc. are taken 1:1 from [9] to ensure unbiased selection. The hyperparameters for unlearning methods are taken from [9] and [25]. For ETD and REM each mask has 20% of the “memorization” partition active during training. The “memorization” partition size is the leftover capacity from “generalization” part to 100% capacity model. For a sensitivity analysis of “generalization” part and mask sizes please refer to [18]. Training time on one A100 (40GB) GPU for one seed for the reported results on ResNet-9 and ViT across training types, discovery rates, regularity levels, any UL method hyperparameters is ca. one week depending on the frequency of logging accuracy results.
Table 3: (Hyper)parameter overview. The UL learning rate for ResNet-9 is the final learning rate of the SGD scheduler during training as set by [9]. We use the same 1/5 fraction from starting learning rate to unlearning learning rate for ViT.
@@ A.2 Model and UL Method Parameters
A.3 Further Ablation Step 3.2
Fig. 6 shows that the barrier step (3.2 in Alg 1) adds (i) stability to the results without the performance jumps observed without the barrier step (ii) and leads to better performance at most discovery rates and regularity levels. The few instances where REM without the barrier step outperforms the default REM indicate that adding the barrier to the loss causes some performance loss which is more pronounced at lower discovery rates when the loss is noisier due to the smaller sample size.
@@
A.4 Model Capacity and Additional ETD Insights
@@ A.5 Detailed Results
Table 4: Aggregate results of methods on CIFAR10 with ResNet-9 (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Error reflects ±$\pm 1$ SEM.
@@
Table 5: Aggregate results of methods on SVHN with ViT (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Values are to the best method for each discovery rate - regularity pairing (best by Utility $\times$ Healed). Potion OOM and thus not reported. Error reflects ±$\pm 1$ SEM.
@@
A.6 Limitations
@@
We highlight the key limitations and assumptions in the main paper. On the unlearning side this is the assumption of having access to the full retain set at unlearning time as is common in the literature. We leave studies of more restrictive settings to future work. The main limitation of REM that harms performance is its masking strategy as highlighted in the paper. We also show the potential that can be unlocked by improved masking with REM (IDEAL) that uses information that is not available in practice to reflect perfect masking. Due to imperfect masking, REM will not perform to its maximum potential in practice, which we call out and provide inspiration for future work to address this. We also note that the removal step can be improved by using future advancements in unlearning methods that do not rely on a retain set to replace our implementation of NPO for this step. Depending on the used method for removal, the limitations of the chosen method will be inherited by REM. We show that REM is robust across tasks, models, optimizers and datasets. Prior UL literature has shown that unlearning methods show stable results across varying dataset sizes in high regularity tasks [25] as well as lower regularity tasks of the same nature (e.g. vision classifiers). REM’s runtime is dependent on the chosen number of epochs and comparable to an epoch of normal training as step 2 of REM uses the much smaller forget set and is negligible in comparison to the backpropagation step on the full training data.

### v2

@@ 7 Discussion and conclusion
We presented a taxonomy for corrupted data unlearning tasks along two dimensions: discovery rate and regularity, and showed that no prior algorithm succeeds across the entire space. We then proposed REM, the first unlearning method that performs strongly across this 2D space of unlearning tasks. REM redirects corrupted data to a dedicated model part that is dropped or deactivated after unlearning to remove the influence of corrupted data from the model. A limitation of REM is that, as in ETD, its masks are binary. Future work could consider softer masking strategies that may allow for better self-organization within $\theta_{o_{2}}$, leading to masks of (even undiscovered) corrupted data having greater overlap with one another. Such better masking strategies may lead to bridging the performance gap with REM (IDEAL) in Fig 3. However, in its current version, REM already makes important strides forward as the first universal method that is strong across our task space. We believe that our 2D framework is an important step forward in better understanding when different families of unlearning algorithms fail, offering useful tools and vocabulary for explaining their behaviours. We hope that future work expands our framework through identifying other key dimensions, builds upon it, for instance by translating its tasks to other modalities, and leverages it for investigating the behaviours of algorithms for other types of unlearning applications.
@@ Appendix A Appendix
A.1 Prior work mapping in our 2D space
The following works are representative for the highlighted areas in Fig. 1 (a).
  • A
    Maini et al. (2023)
  • B
    Schoepf et al. (2024b)
  • C
    Chundawat et al. (2023); Kurmanji et al. (2023)
  • D
    Goel et al. (2024)
  • E
    Zhao et al. (2024); Foster et al. (2024); Chundawat et al. (2023); Kurmanji et al. (2023)
  • F
    Chundawat et al. (2023); Kurmanji et al. (2023)
A.2 Healing Metric
@@
A.3 Model and UL Method Parameters
We perform unlearning with a limited hyperparameter sweep as the best values are already known for high and lower regularity tasks and varying discovery rates from Goel et al. (2024) (poison trigger and interclass confusion tasks). We report the selected hyperparameters in Tab. 3. The ViT architecture and setup for SVHN is adapted from Torkzadehmahani et al. (2024), only changing the optimizer to Adam and adapting the learning rate and batch size accordingly. The ResNet-9 setup as well as the poison trigger and interclass confusion tasks are used from Goel et al. (2024). The random label unlearning problem is adapted from Maini et al. (2023). The REM parameter $\beta$ is kept at 1 to avoid adding extra complexity to the method by introducing weighting.. Training data splits, choice of which sample IDs to manipulate etc. are taken 1:1 from Goel et al. (2024) to ensure unbiased selection. The hyperparameters for unlearning methods are taken from Goel et al. (2024) and Schoepf et al. (2024b). For ETD and REM each mask has 20% of the “memorization” partition active during training. The “memorization” partition size is the leftover capacity from “generalization” part to 100% capacity model. For a sensitivity analysis of “generalization” part and mask sizes please refer to Maini et al. (2023). From our experiments that are seeded with a global seed (affecting masks), we observe that REM is not sensitive to the random seed of the mask. Given the three seeds (0, 1, 2), we get the following healing and utility in the ResNet + CIFAR10 random label unlearning setting for 50% discovery rate (representative halfway point of discovery): Healing [0.772, 0.774, 0.752,], Utility [0.891, 0.888, 0.896]. Results reported for random label unlearning, as this scenario has the least regularity and highest randomness. Training time on one A100 (40GB) GPU for one seed for the reported results on ResNet-9 and ViT across training types, discovery rates, regularity levels, any UL method hyperparameters is ca. one week depending on the frequency of logging accuracy results.
Regarding generalisability across architectures, our experimental results show that REM works across architectures with Transformers and ResNet. The implementation in the ViT is agnostic to vision tasks as we place the REM expansions before/after the FF layers in each transformer block. This setup is applicable to transformer models across modalities but lies outside the scope of this paper. Ghosal et al. (2025) also shows that the ETD approach can be scaled to LLMs for memorization prevention during training (not an unlearning application), indicating that REM is highly likely to succeed at larger scales.
Table 3: (Hyper)parameter overview. The UL learning rate for ResNet-9 is the final learning rate of the SGD scheduler during training as set by Goel et al. (2024). We use the same 1/5 fraction from starting learning rate to unlearning learning rate for ViT.
@@ A.3 Model and UL Method Parameters
A.4 REM Ablation
Extending upon the ablations in Table 2, Fig. 7 shows that the barrier step (3.2 in Alg 1) adds (i) stability to the results without the performance jumps observed without the barrier step (ii) and leads to better performance at most discovery rates and regularity levels. The few instances where REM without the barrier step outperforms the default REM indicate that adding the barrier to the loss causes some performance loss which is more pronounced at lower discovery rates when the loss is noisier due to the smaller sample size.
@@ A.4 REM Ablation
A.5 Model Capacity and Additional ETD Insights
@@ A.6 Detailed Results
Table 4: Aggregate results of methods on CIFAR10 with ResNet-9 (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Error reflects $\pm 1$ SEM.
@@
Table 5: Aggregate results of methods on SVHN with ViT (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Values are to the best method for each discovery rate - regularity pairing (best by Utility $\times$ Healed). Potion OOM and thus not reported. Error reflects $\pm 1$ SEM. No UL and Ascent (best performing prior method) with 100% capacity are reported for reference to the 50% capacity models to show that the observations from ResNet experiments on capacity levels are applicable too.
@@
100 No UL 59.30 $\pm$ 2.43 85.83 $\pm$ 0.56 50.94 $\pm$ 2.92
A.7 Additional experiments with different data, corruption counts, and capacity levels
We train the same ResNet as for CIFAR10 with the same learning rate etc. for the same 40 epochs on Imagenette [^1] https://github.com/fastai/imagenette (see Tab. 6 & 7) but with 100 corruptions (to add a different size for variety). The new experiments again show REM as the top performing method, even with a 90/10 capacity split (Tab. 7). The gap between methods is smaller due to only 100 corruptions having lower influence via reintroduction of corruptions in partial discovery settings (i.e. methods fail later in the high regularity setting). Tables 8 and 9 show the same experiments as in the main body of the paper but with ETD/REM masks that are 1/4 the size of the main body experiments (20% vs 5%) to show that results hold across different sizes and do not vary greatly.
Table 6: Aggregate results of methods on Imagenette with ResNet (50% capacity each, no ETD pretraining) and 100 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Values are to the best method for each discovery rate - regularity pairing (best by Utility $\times$ Healed). Error reflects $\pm 1$ SEM.
UL type Healed Utility Utility $\times$ Healed
REM $76.03\pm 2.61$ $80.32\pm 0.28$ $60.98\pm 2.02$
SCRUB $67.63\pm 3.63$ $80.66\pm 0.17$ $54.52\pm 2.89$
Retrained $65.93\pm 3.04$ $78.17\pm 0.27$ $51.49\pm 2.32$
Potion $56.40\pm 4.84$ $72.44\pm 2.50$ $42.61\pm 3.83$
BadT $47.90\pm 3.18$ $79.64\pm 0.31$ $38.09\pm 2.50$
Ascent $51.37\pm 2.67$ $75.38\pm 1.38$ $38.05\pm 1.66$
No UL $31.00\pm 8.08$ $79.73\pm 1.16$ $24.69\pm 6.38$
Table 7: Aggregate results of methods on Imagenette with ResNet (90% capacity each, with ETD pretraining) and 100 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Values are to the best method for each discovery rate - regularity pairing (best by Utility $\times$ Healed). Error reflects $\pm 1$ SEM.
UL type Healed Utility Utility $\times$ Healed
REM $77.33\pm 2.72$ $79.51\pm 0.21$ $61.41\pm 2.09$
SCRUB $68.83\pm 3.90$ $79.96\pm 0.19$ $55.02\pm 3.07$
Ascent $66.97\pm 2.01$ $75.17\pm 0.73$ $50.21\pm 1.44$
Retrained $64.67\pm 2.93$ $77.54\pm 0.21$ $50.14\pm 2.26$
Potion $64.77\pm 3.34$ $69.90\pm 1.64$ $46.28\pm 3.16$
BadT $55.57\pm 2.49$ $78.59\pm 0.13$ $43.62\pm 1.92$
No UL $45.33\pm 8.25$ $78.50\pm 0.44$ $35.53\pm 6.30$
Table 8: Aggregate results of methods using smaller masks (5% active vs 20% in main experiments) on CIFAR10 with ResNet-9 (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Error reflects $\pm 1$ SEM.
UL type Healed Utility Utility $\times$ Healed
REM 85.37 $\pm$ 1.77 89.56 $\pm$ 0.26 76.38 $\pm$ 1.48
Potion 78.54 $\pm$ 2.45 85.84 $\pm$ 0.68 67.68 $\pm$ 2.41
Ascent 75.72 $\pm$ 3.01 88.75 $\pm$ 0.47 66.81 $\pm$ 2.38
NPO 75.88 $\pm$ 3.25 88.03 $\pm$ 0.53 66.33 $\pm$ 2.54
BadT 64.64 $\pm$ 2.74 89.28 $\pm$ 0.27 57.57 $\pm$ 2.32
Retrained 59.86 $\pm$ 4.94 88.79 $\pm$ 0.24 53.21 $\pm$ 4.39
SCRUB 59.04 $\pm$ 4.92 89.99 $\pm$ 0.25 53.02 $\pm$ 4.36
No UL 51.87 $\pm$ 19.89 89.47 $\pm$ 0.93 46.31 $\pm$ 17.54
Table 9: Aggregate results of methods using smaller masks (5% active vs 20% in main experiments) on SVHN with ViT (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed. Error reflects $\pm 1$ SEM.
UL type Healed Utility Utility $\times$ Healed
REM 81.19 $\pm$ 1.16 85.38 $\pm$ 0.10 69.33 $\pm$ 1.01
Ascent 77.47 $\pm$ 1.80 81.41 $\pm$ 0.85 63.49 $\pm$ 2.03
SCRUB 63.26 $\pm$ 6.28 86.21 $\pm$ 0.09 54.61 $\pm$ 5.43
BadT 61.38 $\pm$ 5.69 83.96 $\pm$ 0.14 51.61 $\pm$ 4.80
No UL 59.90 $\pm$ 24.29 84.40 $\pm$ 0.45 50.64 $\pm$ 20.55
NPO 59.58 $\pm$ 6.42 84.40 $\pm$ 0.12 50.37 $\pm$ 5.43
Retrained 58.81 $\pm$ 4.91 76.32 $\pm$ 0.24 44.98 $\pm$ 3.75
A.8 Compute time and Memory
We report unlearning time as relative numbers compared to initial training for the CIFAR10 runs with a 50% capacity model (average runtimes). Initial training 100%, Bad Teacher 24.7%, SCRUB 14.0%, REM 10.7%, Potion 9.1%, NPO 0.2%, Ascent 0.2%. A key consideration in our benchmark is that no prior method, not even retraining from scratch, can address the 2D space of unlearning tasks presented. This is different to privacy focused unlearning, where retraining from scratch is the gold standard. Therefore, being more efficient is a necessity in privacy focused unlearning - otherwise retraining is better. Given that this is not the case for corrupted data, time is not essential until multiple methods are able to address the 2D space (REM is fast nonetheless compared to training from scratch, SCRUB, etc. and only misses out to methods that do not perform any repair steps). Regarding memory requirements, for REM this depends on the size of $\theta_{o_{2}}$ to keep the expanded model in memory. Some methods such as Potion have higher peak memory usage due to expensive parameter importance computations (see OOM problem with ViT), while others are more efficient due to no $\theta_{o_{2}}$ as in Ascent, or are harder to compare due to multi-model student-teacher setups as in SCRUB and Bad Teacher.
A.9 Limitations
@@
We highlight the key limitations and assumptions in the main paper. On the unlearning side this is the assumption of having access to the full retain set at unlearning time as is common in the literature. We leave studies of more restrictive settings to future work. The main limitation of REM that harms performance is its masking strategy as highlighted in the paper. We also show the potential that can be unlocked by improved masking with REM (IDEAL) that uses information that is not available in practice to reflect perfect masking. Due to imperfect masking, REM will not perform to its maximum potential in practice, which we call out and provide inspiration for future work to address this. We also note that the removal step can be improved by using future advancements in unlearning methods that do not rely on a retain set to replace our implementation of NPO for this step. Depending on the used method for removal, the limitations of the chosen method will be inherited by REM. We show that REM is robust across tasks, models, optimizers and datasets. Prior UL literature has shown that unlearning methods show stable results across varying dataset sizes in high regularity tasks Schoepf et al. (2024b) as well as lower regularity tasks of the same nature (e.g. vision classifiers). REM’s runtime is dependent on the chosen number of epochs and comparable to an epoch of normal training as step 2 of REM uses the much smaller forget set and is negligible in comparison to the backpropagation step on the full training data.